OpenStudy (anonymous):

i dont understand about the square root of -144 idk if its 12 or negative 12

OpenStudy (anonymous):

12

OpenStudy (anonymous):

how is it 12 when its a negative?

OpenStudy (anonymous):

my teacher just told me you will not have a negative square root

OpenStudy (ryan123345):

You cant have a negative square root

OpenStudy (ryan123345):

ex: the square root of 9 is 3 because 3x3=9 right?

OpenStudy (anonymous):

so when you square root a negative number it will really be positive

OpenStudy (ryan123345):

if you want to find the square root of -9 what two numbers are the same and multiply to -9? nothing.

OpenStudy (ryan123345):

When you find the square root of something you are asking yourself what two numbers multiply to that but they have to be the same.

OpenStudy (ryan123345):

if you are finding the square root of 16 it would be 4 but if you are finding the square root of -16 you cant multiply two numbers that are the same to give you -16. -4 times -4 = 16.

mathslover (mathslover):

o.O I think there is misconception among here. See, I can write \(\large{\sqrt{-144}}\) as \(\large{\sqrt{-1} \times \sqrt{144}}\) Right?

mathslover (mathslover):

have you learnt about iota ?

mathslover (mathslover):

iota = \(\sqrt{-1}\) therefore we have the value of sqrt of -144 as 12 * iota

mathslover (mathslover):

\(\large{\sqrt{-144} = \sqrt{144} * \sqrt{-1} = 12 * i}\) i is iota here

OpenStudy (anonymous):

no i havent learned any iota